OpenStudy (anonymous):

Write an equation expressing inverse variation for each exercise below. 1. y varies inversely as x, and y = 25 when x = 3 2. y varies inversely as x, and y = 7 when x = 12

OpenStudy (cwtan):

\(\Huge y\alpha \frac {1}{x}\) y=\(\Huge\frac{k}{x}\) where k is constant 25=\(\Huge \frac {k}{3}\) k=75 so it is y=\(\Huge \frac{75}{x}\) try the 2nd part urself